# GeoPF: Infusing Geometry into Potential Fields for Reactive Planning in Non-trivial Environments

### v1

@@ Abstract
Reactive intelligence remains one of the cornerstones of versatile robotics operating in cluttered, dynamic, and human-centred environments. Among reactive approaches, potential fields (PF) continue to be widely adopted due to their simplicity and real-time applicability. However, existing PF methods typically oversimplify environmental representations by relying on isotropic, point- or sphere-based obstacle approximations. In human-centred settings, this simplification results in overly conservative paths, cumbersome tuning, and computational overhead—even breaking real-time requirements. In response, we propose the Geometric Potential Field (GeoPF), a reactive motion-planning framework that explicitly infuses geometric primitives—points, lines, planes, cubes, and cylinders—into real-time planning. By leveraging precise closed-form distance functions, GeoPF significantly reduces computational complexity and parameter tuning effort. Extensive quantitative analyses consistently show GeoPF’s higher success rates, reduced tuning complexity (a single parameter set across experiments), and substantially lower computational costs (up to 2 orders of magnitude) compared to traditional PF methods. Real-world experiments further validate GeoPF’s robustness and practical ease of deployment. GeoPF provides a fresh perspective on reactive planning problems driving geometric-aware temporal motion generation, enabling flexible and low-latency motion planning suitable for modern robotic applications.
@@ I Introduction
Robotics has expanded beyond structured, well-defined factory floors into (i) unstructured, (ii) cluttered, and (iii) dynamic human-centred spaces, as illustrated in Fig. 1. As robots increasingly share workspaces with humans, the need for reactive intelligent navigation solutions becomes paramount. Yet, despite significant advances in motion planning [1], autonomous reactive behavior in settings (i)-(iii) remains an open problem [2, 3, 4]. The core challenge lies in balancing two competing demands: efficiently modeling unstructured and cluttered environments (i)-(ii) while ensuring real-time reactivity to dynamic, unforeseen changes (iii) [5, 6]. This challenge is exemplified in Fig. 1, where a robot must navigate a constrained environment to pour tea, set a breakfast table, and dynamically respond to a human moving objects in real-time. Achieving autonomy and reactive intelligence in such scenarios requires a structured and stable environment representation—one that captures both spatial and temporal information—with a mapping that seamlessly integrates this knowledge into reactive motion generation alongside control constraints and real-time actions. This work addresses this open challenge by providing a novel solution for reactive intelligence, the centerpiece of truly versatile robotic systems operating in human-centered non-trivial environments [7].
Traditional monolithic Sense-Plan-Act (SPA) architectures divide the system into sequential sensing, planning, and action phases [1]. This strong modularization enables robots to function effectively in structured, pre-modeled environments [8, 9, 10]. The key advantage of this paradigm lies in breaking down complex tasks into manageable subproblems, making planning more intuitive and computationally feasible [3]. Most successful approaches including graph-based planners, such as Sparse Roadmaps [11] and Probabilistic Roadmaps (PRMs)[12], construct a free configuration space (C-space) representation and connect edges using local planners, often trading optimality for computational efficiency (e.g., A* [13, 14]). In contrast, tree-based single-query planners, such as Rapidly-exploring Random Trees (RRTs), iteratively expand toward the goal while verifying collision-free paths [15]. Despite their effectiveness in structured settings, these methods lack real-time adaptability. Indeed, while SPA methods excel in handling unstructured (i) and cluttered (ii) scenarios, they fail to adapt efficiently to dynamic (iii) changes [3, 4]. Their inherent sequential nature hinders real-time feedback, making them ill-suited for scenarios where reactive behavior and continuous adaptation to environmental changes are critical [16].
@@
In contrast, vector-field-based motion generation, often termed control-based planning, enables rapid adaptation to environmental changes [17], offering an efficient solution to reactivity (iii). These methods often use gradient-based techniques to guide motion generation [18, 19, 20], ensuring real-time responsiveness. While early implementations suffered from local minima, e.g., artificial potential fields (APFs), these have been mitigated by force-based strategies such as circular fields [21, 22, 23, 4] and magnetic fields [24, 25, 26]. Furthermore, hybrid approaches have attempted to integrate C-space planners with vector-field strategies, leveraging an initial trajectory while benefiting from the reactivity of field-based methods [27, 28]. Learning-based approaches have also been explored to enhance field-based planners [29]. However, all these methods still introduce additional computational complexity without resolving the core bottleneck of field-based methods—their reliance on simplified environment representation.
@@
Despite their computational efficiency compared to SPA, these methods rely on oversimplified workspace representations. Inspired by physics, they often model their interaction via particle-based force dynamics, which assumes point/spherical obstacle representations. While intuitive, this approach poorly represents complex objects like shelves, forcing a trade-off: large repulsive spheres result in overly conservative planning, whereas multiple smaller spheres, ellipsoids, or point clouds lead to excessive distance computations, making real-time execution impractical. Conversely, sparsely selecting sample points causes force discontinuities, complex switching strategies, and short-sighted environment representations, ultimately leading to suboptimal or infeasible trajectories that are highly sensitive to tuning parameters.
@@
In summary, despite the promise of field-based planners in providing intuitive and reactive behaviours, their application to real-world robotic manipulation remains constrained by three fundamental issues: (i) high computational costs associated with representing large, non-isotropic vector fields; (ii) limited performance due to the reliance on isotropic and point-based distance fields; and (iii) the complex and impractical tuning requirements necessary to handle real-world scenarios effectively. These shortcomings are inherently tied to the environment representation upon which existing field-based methods are built. In most cases, obstacles are represented using point clouds [23, 4], or through spheres or octomaps [30, 17, 31] leading to challenges in handling non-spherical objects and non-isotropic vector fields. This results in either excessive computational overhead or complex non-intuitive and cumbersome tuning to handle sparse and changing obstacle representation—both cases leading to poor or even infeasible trajectory generation.
@@
In this work, we take a fundamentally different approach to reactive planning by casting geometric primitives for improved environment representation into force generation. Instead of focusing on incremental improvements to field-based methods—refining point-based techniques, improving tuning, or modifying isotropic fields—we propose a geometric-based representation that explicitly incorporates the spatial properties of objects into the temporal motion generation process. Our key contributions include: (C1) A novel geometric-aware environment representation for collision avoidance, leveraging fundamental geometric primitives (points, lines, planes, cubes, and cylinders) to define more structured avoidance strategies; (C2) The design of a reactive planning strategy, the Geometric Potential Field (GeoPF), based on geometric distances, which surpasses existing methods in computational efficiency, trajectory feasibility, and robustness while requiring minimal tuning; (C3) Extensive quantitative comparisons and real-world experiments, demonstrating that GeoPF achieves orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates and superior trajectory quality.
@@ II GeoPF: System Design and Overview
This work is about reactive planning in non-trivial human-centric environments. Although straightforward extensions are possible for the whole-body [32], our current investigation focuses on the end-effector (EE) $\underline{\bm{x}}{\in}SE(3)$. Let’s consider an initial and a goal pose, respectively $\bm{\underline{x}}^{0},\bm{\underline{x}}^{g}$. Both rotational and translational components are modelled through unit-dual quaternion parametrization, and while the orientation tracking is achieved through screw interpolation [33, 34], in this work, we are interested in the EE cartesian trajectory critical for collision avoidance. This is modelled as a point mass ${\bm{\underline{p}}^{R}}{\in}\mathbb{R}^{3}$.
@@
$\mathcal{T}_{\varnothing}$ Unfeasible Workspace
$\mathcal{T}^{feas}$ Feasible Workspace
@@
$\bm{\underline{\mu}}$ Unit direction in Cartesian Space
$\bm{\underline{f}}$ Force unit direction
@@
$r^{Geo}$ Radius of sphere, circle, cylinder
$\bm{\underline{p}}_{O}^{Geo}$ Center of sphere, circle, cylinder
@@ III From Geometry to Force Generation
One of the key contributions of this work is a geometric-based environment representation that explicitly encodes the spatial properties of objects and uses this geometry to drive temporal motion generation. Rather than relying solely on point clouds or large sets of spheres for obstacle modeling, we cast geometric primitives—including spheres, lines, planes, cubes, and cylinders—into repulsive forces within a vector field.
@@ III-C Plane-based Potential Field
The perpendicular foot lies inside the plane if and only if all indicators $\bm{\underline{c}}_{i}$ point in the same direction, i.e., $\bm{\underline{c}}_{1}{=}\bm{\underline{c}}_{2}{=}\bm{\underline{c}}_{3}{=}\bm{\underline{c}}_{4}$, as seen in Fig. 4(a). In this case, the perpendicular foot lies inside the plane boundaries $\bm{\underline{p}}^{P}_{\bot R}\in\mathcal{T}^{P}$, and the repulsive force direction is orthogonal to the plane and aligned to (8), that is, $\bm{\underline{f}}_{o}^{P}=\bm{\underline{\mu}}^{P}=\bm{\underline{c}}_{i}$.
@@
If the line from robot to goal intersects the plane interior, we project a repulsive force parallel to the plane’s surface (toward the nearest plane edge) rather than purely along the normal. This prevents the attractive force from canceling out the plane’s repulsion, allowing the robot to “slide” around or above the plane.
@@
By defining the nearest edge direction $\bm{\underline{\mu}}^{P}_{e}$, the parallel direction is given by $\bm{\underline{\mu}}^{P}_{corr}=\bm{\underline{\mu}}^{P}\times\bm{\underline{\mu}}^{P}_{e}$. To avoid crossing toward the farthest edge, see Fig. 5(b), we evaluate the direction from the robot to the nearest edge, denoted as $\bm{\underline{\mu}}^{PR}_{e}$, which corresponds to the opposite direction of the line-to-robot vector, Section III-B. By computing $s=\text{sign}(<\bm{\underline{\mu}}^{P}_{corr},\bm{\underline{\mu}}^{PR}_{e}>)$, we can determine whether the parallel direction toward the nearest edge or its opposite. When $s=1$, the direction is correctly toward the nearest edge ($\alpha<\frac{\pi}{2}$), whereas $s=-1$ indicates the farthest edge ($\alpha>\frac{\pi}{2}$), see Fig. 5. Accordingly, the corrected force is
@@
orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates and superior trajectory quality.orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates and superior trajectory quality.
@@ III-D Geometric Cube-Based Potential Field
Figure 6: Cubic force relies on the closest surface. (a) When robot closest to surface, the surface plane provides the force (Sec. III-C. (b) When robot closest to an edge, either surface can be chosen to calculate the cubic force.
@@ III-E Cylinder-Based Potential Field
Cylinders commonly arise from objects such as cups, bottles, cans, pipes, or cylindrical machinery components. We define a cylinder by its central axis (stemming from two vertices $\bm{\underline{p}}^{Y}_{O1},\bm{\underline{p}}^{Y}_{O2}{\in}\mathbb{R}^{3}$) and the radius $r^{Y}$. We explore the curved surface surrounding the cylinder, and two circular-cross section planes centred at the vertices. The normal direction of the curved surface is the radial direction from the central axis extending outward, Fig 7. The normal directions of the circular surfaces are perpendicular to the plane and point outside the cylinder, then we directly apply the force along to the axis direction $\bm{\underline{f}}^{Y}_{X}$.
@@ IV Quantitative Analysis and Experiments
In this section, we present an extensive evaluation of our proposed GeoPF in both simulated and real-world settings. Our goal is twofold: (i) To quantitatively assess performance [^1] Key performance metrics include (i) success rate—whether the robot reaches the goal without collision; (ii) computational time—the total time (or average per-step time) for generating a full trajectory; (iii)-(iv) path quality—including path length and minimum distance to obstacles. against baseline reactive planners based on state-of-the-art field-based methods across a variety of randomized scenarios (Subsection IV-A), and to demonstrate GeoPF’s feasibility and reliability in real robotic experiments with a Franka Emika arm (Subsection IV-B).
To ensure fairness, all experiments were run on the same hardware. [^2] AMD Ryzen 7 5800H, 32GB DDR4-3200 under Ubuntu 20.04, gcc/c++ 9.4. The randomization seeds, number of obstacles, and various obstacle positions for each scenario vary to thoroughly test each method’s robustness.

### v2

@@ Abstract
Reactive intelligence remains one of the cornerstones of versatile robotics operating in cluttered, dynamic, and human-centred environments. Among reactive approaches, potential fields (PF) continue to be widely adopted due to their simplicity and real-time applicability. However, existing PF methods typically oversimplify environmental representations by relying on isotropic, point- or sphere-based obstacle approximations. In human-centred settings, this simplification results in overly conservative paths, cumbersome tuning, and computational overhead—even breaking real-time requirements. In response, we propose the Geometric Potential Field (GeoPF), a reactive motion-planning framework that explicitly infuses geometric primitives—points, lines, planes, cubes, and cylinders—their structure and spatial relationship in modulating the real-time repulsive response. Extensive quantitative analyses consistently show GeoPF’s higher success rates, reduced tuning complexity (a single parameter set across experiments), and substantially lower computational costs (up to 2 orders of magnitude) compared to traditional PF methods. Real-world experiments further validate GeoPF’s reliability, robustness, and practical ease of deployment, as well as its scalability to whole-body avoidance. GeoPF provides a fresh perspective on reactive planning problems driving geometric-aware temporal motion generation, enabling flexible and low-latency motion planning suitable for modern robotic applications.
@@ I Introduction
Robotics has expanded beyond structured, well-defined factory floors into (i) unstructured, (ii) cluttered, and (iii) dynamic human-centred spaces, increasing the demand for reactive intelligent navigation, Fig. 1. Yet, despite significant advances in motion planning [1], autonomous reactive behavior remains an open problem [2, 3]. The core challenge lies in addressing clutter while providing structured environment representations that enable real-time reactivity to dynamic, unforeseen changes [4, 5]. This challenge is exemplified in Fig. 1. A robot must navigate a constrained environment to pour tea, set a table, and dynamically respond to a human moving objects in real-time. This requires a structured environment representation—one capable of seamlessly integrating geometric, spatial, and temporal information into real-time reactive planning. This work addresses precisely this open challenge by introducing a novel paradigm that explicitly leverages shape-awareness, geometric structure, and spatial relationship into repulsive strategies beyond single-gradient avoidance—enabling reactive intelligence, the centerpiece of truly versatile systems operating in human-centered non-trivial environments [6].
Traditional monolithic Sense-Plan-Act (SPA) architectures divide the system into sequential sensing, planning, and action phases [1]. This strong modularization enables robots to function effectively in structured, pre-modeled environments [7, 8, 9]. The key advantage lies in breaking down complex tasks into manageable subproblems, making planning more intuitive and computationally feasible [2]. Most successful approaches including graph-based planners construct a free configuration space (C-space) representation and connect edges using local planners [1, 10, 9]. Instead, tree-based single-query planners, as Rapidly-exploring Random Trees (RRTs), iteratively expand toward the goal through collision-free paths [10]. Despite their effectiveness in structured settings, they lack real-time adaptability. Indeed, SPA methods excel in handling unstructured (i) and cluttered (ii) scenarios, but they fail to adapt efficiently to dynamic (iii) changes [2, 3]. Their inherent sequential nature hinders real-time feedback, making them ill-suited for scenarios where reactive behavior and continuous adaptation to environmental changes are critical [11].
@@
In contrast, vector-field-based motion generation, or reactive planning, efficiently addresses the real-time responsiveness challenge (iii). These methods use gradient-based techniques to guide motion generation [12, 13, 14]. While early implementations suffered from local minima, these have been mitigated by strategies such as circular fields (CFs) [15, 16, 17, 3] and magnetic fields [18, 19, 20], which incorporated the sphere geometry to generate tangential repulsive forces to complement the standard repulsive gradient.
While mitigating traditional drawbacks (e.g., local minima), these methods fundamentally rely on point or sphere-based obstacle representations. This contrasts with modern simulators, engines and perception algorithms for collision detection (see e.g. MuJoCo [21], GJK [22], and FCL [23]) that compute (analytically or iteratively) contact-point distances and normals from richer geometric information. Likewise, voxel- and Signed Distance Function (SDF)-based methods [24, 25, 26] (as well as neural SDFs variants [27]) compute geometry-aware closest-point smooth gradients for collision avoidance. Notwithstanding, during motion generation, all the aforementioned techniques reduce the obstacle interaction to a scalar distance with a repulsive to the closest point alone [28] or local gradients [26] with respect to the closest point. That is, the rich information about the obstacle geometry, structure and spatial configuration is only used to compute the closest point and then neglected during motion generation.
@@
In other words, despite their computational efficiency compared to SPA, existing reactive planners remain fundamentally constrained to a single-point or single-gradient, towards the closest point, interaction—disregarding the richer geometry and structural information inherent to the object. While intuitive, this particle-based approach provides poor information for the motion generation strategy. When considering point-clouds or point-based geometry, these methods lead to either large repulsive spheres yielding overly conservative plans, or dense sampling, which increases computational load, making real-time execution impractical. Conversely, sparse sampling causes force discontinuities, switching strategies, and short-sighted environment representations, ultimately leading to suboptimal or infeasible trajectories that are highly sensitive to tuning parameters.
@@
In summary, despite the promise of field-based planners in providing intuitive and reactive behaviours, their application remains constrained by three fundamental issues: (i) high computational costs associated with representing large, non-isotropic vector fields; (ii) limited performance due to the reliance on isotropic,or closest-point-based distance fields and gradients; and (iii) the complex and impractical tuning requirements necessary to handle real-world scenarios effectively. These shortcomings are inherently tied to the environment representation upon which existing field-based methods are designed. In most cases, repulsive forces are build upon point clouds based distances [17, 3], or through spheres or octomaps [29, 30], leading to challenges in handling non-spherical objects and non-isotropic vector fields. While geometric-aware distance computation methods, e.g., [21, 22, 23, 24, 25, 26] help reduce the computational load (i), the resulting reactive planners still ultimately reduce obstacles interactions to a single pointwise gradient, using the nearest-distance vector irrespective of whether the closest-contact lies on a face, an edge, or a corner. As a result, planners must rely on local contact information only, which limits geometric expressiveness and leads to non-intuitive force profiles. This lack of structural awareness often requires non-intuitive and cumbersome tuning leading to poor or even infeasible trajectory generation.
@@
The key novelty of this work lies in exploiting the rich structure of geometric primitives to directly shape the reactive vector field. Rather than relying solely on the closest-point distance, GeoPF leverages the structural representation of each primitive. These features—promptly available in our method and other geometric modeling pipelines—are most often overlooked during the planning stage. Similarly, in motion optimization frameworks, focus lies not on the shape itself but on how far the robot is from violating a closest-point contact-based threshold. Thus, local point-based gradients dictate the collision logic instead of distinguishing between the robot approaching a face, corner, or a cylindrical wall. To the best of our knowledge, such information indeed has never been used during reactive planning and field generation.
Our key contributions include: (C1) A novel geometric-aware planning paradigm, revising and generalizing primitives, leveraging their topological and structural properties to actively modulate the repulsive field beyond simple distance-based computations; (C2) An analytical, closed-form vector-field planner explicitly tailored for common primitives (points, lines, planes, cubes, and cylinders), directly instantiating C1; (C3) Comprehensive quantitative analysis highlighting GeoPF’s higher success rates and orders-of-magnitude computational improvements over state-of-the-art field-based planners, accompanied by real-world experiments — including both whole-body avoidance and poor/coarse primitive shaping showcasing its scalability and robustness.
@@ II GeoPF: System Design and Overview
This work is about reactive planning in non-trivial human-centric environments. $\underline{\bm{x}}{\in}SE(3)$. Let’s consider an initial and a goal pose, respectively $\bm{\underline{x}}^{0},\bm{\underline{x}}^{g}{\in}SE(3)$. Both rotational and translational components are modelled through unit-dual quaternion parametrization, and while the orientation tracking is achieved through screw interpolation [31], in this work, we are interested in the Cartesian trajectory, ${\bm{\underline{p}}^{R}}{\in}\mathbb{R}^{3}$.
@@
$\mathcal{T}^{feas}$, $\mathcal{T}_{\varnothing}$ Feasible and Unfeasible Workspace
@@
$\bm{\underline{\mu}}$, $\bm{\underline{f}}$ Unit and Force Unit direction in Cartesian Space
@@
$r^{Geo}$, $\bm{\underline{p}}_{O}^{Geo}$ Radius and Center of sphere, circle, cylinder
@@ III From Geometry to Force Generation
One of the key contributions of this work is a geometric-based environment representation that explicitly encodes the spatial properties of objects and uses this geometry to drive temporal motion generation. Rather than relying solely on point clouds or large sets of spheres for obstacle modeling, or the closest-point contact-based to the obstacle, we cast primitives—including spheres, lines, planes, cubes, and cylinders—into repulsive forces within a vector field.
@@ III-C Plane-based Potential Field
The perpendicular foot lies inside the plane if and only if all indicators $\bm{\underline{c}}_{i}$ point in the same direction, i.e., $\bm{\underline{c}}_{1}{=}\bm{\underline{c}}_{2}{=}\bm{\underline{c}}_{3}{=}\bm{\underline{c}}_{4}$, Fig. 4(a). In this case, the perpendicular foot lies inside the plane boundaries $\bm{\underline{p}}^{P}_{\bot R}\in\mathcal{T}^{P}$, and the repulsive force direction is orthogonal to the plane and aligned to (8), i.e., $\bm{\underline{f}}_{o}^{P}=\bm{\underline{\mu}}^{P}=\bm{\underline{c}}_{i}$.
@@
If the line from robot to goal intersects the plane interior, we project a repulsive force parallel to the plane’s surface (toward the nearest plane edge) rather than purely along the normal. This prevents the attractor from canceling out the plane’s repulsion, allowing the robot to “slide” around or above the plane.
@@
By defining the nearest edge direction $\bm{\underline{\mu}}^{P}_{e}$, the parallel direction is given by $\bm{\underline{\mu}}^{P}_{corr}=\bm{\underline{\mu}}^{P}\times\bm{\underline{\mu}}^{P}_{e}$. To avoid crossing toward the farthest edge, see Fig. 5(b), we evaluate the direction from the robot to the nearest edge [^1] In rare case of exact symmetry, similarly to [3], a small random additive noise is applied toward one of the primitive’s vertices or edges., denoted as $\bm{\underline{\mu}}^{PR}_{e}$, which corresponds to the opposite direction of the line-to-robot vector, Section III-B. By computing $s=\text{sign}(<\bm{\underline{\mu}}^{P}_{corr},\bm{\underline{\mu}}^{PR}_{e}>)$, we can determine whether the parallel direction toward the nearest edge or its opposite. When $s=1$, the direction is correctly toward the nearest edge ($\alpha<\frac{\pi}{2}$), whereas $s=-1$ indicates the farthest edge ($\alpha>\frac{\pi}{2}$), see Fig. 5. Accordingly, the corrected force is
@@
GeoPF has orders-of-magnitude improvements in computational speed over state-of-the-art reactive field-based planners while maintaining higher success rates.
@@ III-D Geometric Cube-Based Potential Field
Figure 6: Cubic force relies on the closest surface. (a) When robot closest to surface, the surface plane provides the force (Sec. III-C. (b) When robot closest to an edge, either surface can be chosen.
@@ III-E Cylinder-Based Potential Field
Cylinders commonly arise from objects such as cups, bottles, cans, pipes, or cylindrical machinery components. We define a cylinder by its central axis (stemming from two vertices $\bm{\underline{p}}^{Y}_{O1},\bm{\underline{p}}^{Y}_{O2}{\in}\mathbb{R}^{3}$) and the radius $r^{Y}$. We explore the curved surface surrounding the cylinder, and two circular-cross section planes centred at the vertices. The normal direction of the curved surface is the radial direction from the central axis extending outward, Fig 7. The normal of the circular surfaces are perpendicular to the plane and point outside the cylinder, then we directly apply the force along the axis direction $\bm{\underline{f}}^{Y}_{X}$.
@@
((a)) Comparison of GeoPF and CF Spherization in Easy (Left) and Hard (Right) Scenarios. Black lines and circles represent line obstacles and their spherization. In clutter, spherization forces trajectories around all obstacles, while GeoPF navigates narrow spaces between obstacles and successfully reaches the goal.
((b)) Spherization of Line and Plane
((c)) Trajectory in Plane Scenario
Figure 8: Spherization in random Line and Plane Scenes. (b) shows how spherization scales objects from lines to planes, sharply increasing computation time. (c) GeoPF finds feasible paths, CF and PF are trapped in local minima, yielding lower success rates.
@@ IV Quantitative Analysis and Experiments
This section presents an extensive evaluation of GeoPF with two main goals, (i) to quantitatively assess its performance [^2] Key performance include (i) success rate—reaching the goal w/o collision, (ii) computational time; (iii)-(iv) path length and min. distance to obstacles. against state-of-the-art field-based planners across a variety of randomized scenes, and to demonstrate GeoPF’s feasibility, reliability, and robustness in real experiments with a Franka robot. To ensure fairness, all experiments were run on the same hardware, [^3] AMD Ryzen 7 5800H, 32GB DDR4-3200 under Ubuntu 20.04, gcc/c++ 9.4. with randomized seeds, obstacle counts and placements, to thoroughly test robustness.